OpenStudy (anonymous):

Ride a bike 12 miles an hour, drive in the car 36 miles back...round trip was 1 hour. How far was the distance?

OpenStudy (ash2326):

Bike's speed 12 miles/hr Car covers 36 miles/hour both covers equal distance x time taken by bike x/12 hours time taken by car x/36 total time=1 hour x/12+x/36=1 3x+x=36 x=9 Distance is 9 miles

OpenStudy (anonymous):

Yeah correct, 9 miles.